OpenStudy (anonymous):

What the solution to: 6v+7 = -29?

OpenStudy (anonymous):

Step by step. Step 1: Sobtract 7 from both sides. What do you get?

OpenStudy (anonymous):

what is your answer.. ?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

6v+0=-36

OpenStudy (anonymous):

then ? what is the next step to find v :) ?

OpenStudy (anonymous):

Okay. Now divide both sides by 6. That gives you the value of v.

OpenStudy (anonymous):

6/6=1 36/6=6

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

nonono

OpenStudy (anonymous):

No. Erm... see, let's not write 0 in there, it's just 6v=-36 okay?

OpenStudy (anonymous):

now, when I divide both sides by 6, what does that leave me with?

OpenStudy (anonymous):

i'm not great with math...

OpenStudy (anonymous):

6v = - 36 (divide both sides by 6) 6v/6 = - 36/6 v = can you finish this ?

OpenStudy (anonymous):

@Camtron it's not math. It's logic. Try seeing it as logic, instead of numbers and varibles. Okay? The worst thing that can happen is a mistake. Doesn't matter. Mistakes can be corrected.

OpenStudy (anonymous):

all you have to do now is divide...you can do that :)

OpenStudy (anonymous):

step 1